SUPERSONICALLY COOLED MOLECULAR IONS IN A SLIT-JET DISCHARGE: HIGH-RESOLUTION INFRARED SPECTROSCOPY AND TUNNELING DYNAMICS OF HD_{2}O^{+}

\underline{FENG DONG}, DAIRENE UY, SCOTT DAVIS, MARK CHILD, and DAVID J. NESBITT, JILA, University of Colorado, Boulder, CO 80309.

Jet-cooled high-resolution infrared spectra of partially deuterated hydronium ion, HD_{2}O^{+} in the O-H stretch region (ν_{3} band) are obtained for the first time, exploiting the high ion densities, long absorption path lengths, and concentration modulation capabilities of the slit jet discharge spectrometer. Least-square analysis with a Watson asymmetric top Hamiltonian yields rovibrational constants and provides high level tests of ab initio molecular structure predictions. Transitions out of both lower (ν_{3}^{+} ← 0^{+}) and upper (ν_{3}^{−} ← 0^{−}) tunneling levels are observed, as well as transitions across the tunneling gap (ν_{3}^{−} ← 0^{+}). The ν_{3}^{−} ← 0^{+} transitions in HD_{2}O^{+} acquire oscillator strength by loss of C_{3v} symmetry and permit both ground (27.0318(72) cm^{−1}) and excited state (17.7612(54) cm^{−1}) tunneling splittings to be determined to spectroscopic precision from a single rovibrational band. The splittings and band origins calculated with recent high level ab initio 6D potential surface predictions for H_{3}O^{+} and isotopomers ^{a} ^{b} are in very encouraging agreement with the current experimental results.

aX. C. Huang, S. Carter, and J. M. Bowman, J. Chem. Phys. 118 (12), 5431 (2003).

bT. Rajamaki, A. Miani, and L. Halonen, J. Chem. Phys. 118 (24), 10929 (2003).